Janet is going to purchase a new bedroom set from Ashley Furniture. She finds a set that she loves and will cost her $5,550 for all items. She is choosing to do an 8-month deferment payment. How much interest will she have gain onto her cost is she is charged an 22% rate compounded monthly?
1
Expert's answer
2020-09-28T20:05:32-0400

F "=P(1+\frac{r}{100})" n

5500 "=P(1+\frac{22}{100})" .677

P= 4752.56

Interest gain = 5500-4752.56

= 747.44